# From Formal Language Theory to Statistical Learning: Finite Observability of Subregular Languages

### v1

@@ 5.1.1 Strictly Local Languages (SLk)
We generated synthetic datasets for SL3 languages. A set of forbidden 3-grams was defined over a finite alphabet with explicit inclusion of boundary symbols <#\#>, </#/\#>. Positive examples were sampled uniformly from sequences that avoided all forbidden 3-grams, while negative examples were constructed to guarantee the presence of at least one forbidden 3-gram (including those at boundaries). Each string length was sampled randomly within a fixed range (5–15 characters). The feature representation was the full set of boundary-aware 3-grams over the extended alphabet $\Sigma\cup\{\text{<}\#\text{>},\text{<}/\#\text{>}\}$. This guarantees that the deciding predicates are included.
@@ Results.
5.1.2 Strictly Piecewise Languages (SPk)
@@
For SP2 languages, we defined a forbidden set of subsequences of length two (e.g., $(a,c)$, $(b,d)$) over an alphabet of four symbols. Positive examples were strings avoiding these subsequences, while negative examples were constructed to ensure that at least one forbidden subsequence occurred. String lengths were randomly sampled in the range 6–18 characters. The feature representation was the set of all possible symbol pairs (ordered subsequences) over the alphabet. This again ensures that all deciding predicates are explicitly represented in the feature space.
@@ Results.
As with SL3, Figure 1 (B) shows accuracy under a noise sweep, and Figure 1 (E) shows results for a size sweep. In the absence of label noise, classification achieves 100% accuracy, exactly matching theoretical expectations. When noise is introduced, performance degrades smoothly, illustrating the robustness of the separating representation.
5.1.3 Locally Threshold Testable Languages (LTTk)
@@
For the LTT2 condition, we defined membership in terms of thresholded counts of substrings with boundaries. Specifically, positive examples were required to contain at least two instances of the symbol $a$, begin with $a$ at the left boundary, contain at most one occurrence of the 2-gram $bb$, and contain at most one occurrence of $c$ at the right boundary. Positive examples were generated by rejection sampling with targeted repairs to ensure that all threshold constraints were satisfied. Negative examples were produced by minimally breaking one constraint, for example by reducing the number of $a$’s, inserting an additional $bb$, altering the prefix so it no longer begins with $a$, or forcing a terminal $c$. This guaranteed that every negative sequence violated at least one deciding predicate. Feature vectors were constructed from all substrings of length at most two, extended with boundary markers, with each feature indicating whether the count of that substring exceeded a given threshold. These features explicitly encode the deciding predicates of LTT2.
@@ Results.
Figure 1 (C) shows the effect of label noise on classification accuracy, and Figure 1 (F) shows the effect of training size. In the noise-free setting, accuracy reached 100%, confirming the theoretical prediction of perfect separability. As noise increased, accuracy decreased, while larger training sizes quickly restored performance toward the ideal. These results provide empirical evidence that threshold-based constraints in LTT languages are finitely observable and linearly separable when represented with appropriate features.
5.2 Low-Quantile Margins under Label Noise
Figure 2: 1% quantile normalized margin vs. label noise for SL, SP, and LTT languages.
Setup.
To further probe the robustness of linear separability, we measured how the low-quantile normalized margins behave as label noise increases. For each subregular class: Strictly Local (SL), Strictly Piecewise (SP), and Locally Threshold Testable (LTT), we trained logistic regression classifiers on synthetic datasets of size 2,000, where a fraction of the training labels were flipped at random. Test sets of 1,000 sequences per class were kept noise-free. After training, we computed the normalized margin for each test example, defined as $m_{i}\;=\;y_{i}\cdot\frac{w^{\top}x_{i}+b}{\lVert w\rVert_{2}}$ and extracted the 1st percentile ($q_{0.01}$) across the distribution. This quantile analysis highlights the fragility of the worst-classified points, thus serving as a probe of separability under noise.
Results.
Figure 2 shows the 1% quantile margin as a function of label noise. SL (light blue) maintains relatively high margins at low noise levels and exhibits a gradual decline, indicating strong robustness of local constraints. SP (light orange) drops more quickly with noise, reflecting the fact that long-distance constraints are easily disrupted by even a small number of mislabels. LTT (light green) displays consistently negative margins, as its threshold-based definition inherently places a significant portion of examples near the decision boundary; nevertheless, its degradation with noise is moderate and smoother than that of SP.
@@ 6 Experiments on Morphological Data
We represented each affix sequence using two subregular feature classes: Piecewise Testable (PT) predicates, capturing all subsequences up to length $m=2$, and Locally Threshold Testable (LTT) predicates, encoding boundary-aware $k=2$. As in the synthetic experiments, we trained logistic regression classifiers on these feature vectors. Importantly, the classifiers were not constructed directly from the theoretical separating hyperplanes, but were trained from data, demonstrating that linear separability is empirically recoverable by standard learning algorithms.
@@ Results.
On the held-out test set, the classifiers achieved strong performance. Logistic regression with PT+LTT features typically reached accuracy above 85%, with F1 scores 86%. Confusion matrices in Figure 3 showed that more errors occurred in negative items that were close to positive forms in surface structure. To better understand what the classifier has learned from the real morphological data, we inspected the top-weighted predicates in the logistic regression model (Figure 4). Strikingly, the most influential features correspond to linguistically interpretable affix constraints. For instance, the model strongly weights boundary-sensitive predicates such as the presence of -ly or -ness at the right edge, reflecting the well-attested distributional restrictions of these suffixes. Similarly, certain prefix–suffix interactions emerge as highly discriminative, such as restrictions involving dis- or re- in combination with other affixes. These results demonstrate that the classifier is not only achieving high predictive accuracy, but is also recovering linguistically meaningful generalizations that align with theoretical expectations in morphology.
@@ 7 Conclusion
SLkk $k$-gram presence in $\tilde{x}$ forbidden substrings
SPkk $k$-subsequence presence in $\tilde{x}$ forbidden subsequences
LTkk $k$-grams + (k1)(k{-}1) prefix/suffix local testability
PTmm subsequences of length $\leq m$ Simon’s theorem
[…]
@@ A.4 Finite Observability of Standard Subregular Classes
Families like $\mathrm{PT}=\bigcup_{m}\mathrm{PT}_{m}$ or $\mathrm{SL}=\bigcup_{k}\mathrm{SL}_{k}$ are directed unions of finitely observable subclasses. Unless parameters are bounded, no single finite predicate set suffices for the whole union; this does not affect applications where the parameters are fixed by design.

### v2

@@ 5.1.1 Strictly Local Languages ($\mathsf{SL}_{k}$)
We generated synthetic datasets for $\mathsf{SL}_{3}$ languages. A set of forbidden 3-grams was defined over a finite alphabet with explicit inclusion of boundary symbols <#\#>, </#/\#>. Positive examples were sampled uniformly from sequences that avoided all forbidden 3-grams, while negative examples were constructed to guarantee the presence of at least one forbidden 3-gram (including those at boundaries). Each string length was sampled randomly within a fixed range (5–15 characters). The feature representation was the full set of boundary-aware 3-grams over the extended alphabet $\Sigma\cup\{\text{<}\#\text{>},\text{<}/\#\text{>}\}$. This guarantees that the deciding predicates are included.
@@ Results.
5.1.2 Strictly Piecewise Languages ($\mathsf{SP}_{k}$)
@@
For $\mathsf{SP}_{2}$ languages, we defined a forbidden set of subsequences of length two (e.g., $(a,c)$, $(b,d)$) over an alphabet of four symbols. Positive examples were strings avoiding these subsequences, while negative examples were constructed to ensure that at least one forbidden subsequence occurred. String lengths were randomly sampled in the range 6–18 characters. The feature representation was the set of all possible symbol pairs (ordered subsequences) over the alphabet. This again ensures that all deciding predicates are explicitly represented in the feature space.
@@ Results.
As with $\mathsf{SL}_{3}$, Figure 1 (B) shows accuracy under a noise sweep, and Figure 1 (E) shows results for a size sweep. In the absence of label noise, classification achieves 100% accuracy, exactly matching theoretical expectations. When noise is introduced, performance degrades smoothly, illustrating the robustness of the separating representation.
5.1.3 Locally Threshold Testable Languages ($\mathsf{LTT}_{k}$)
@@
For the $\mathsf{LTT}_{2}$ condition, we defined membership in terms of thresholded counts of substrings with boundaries. Specifically, positive examples were required to contain at least two instances of the symbol $a$, begin with $a$ at the left boundary, contain at most one occurrence of the 2-gram $bb$, and contain at most one occurrence of $c$ at the right boundary. Positive examples were generated by rejection sampling with targeted repairs to ensure that all threshold constraints were satisfied. Negative examples were produced by minimally breaking one constraint, for example by reducing the number of $a$’s, inserting an additional $bb$, altering the prefix so it no longer begins with $a$, or forcing a terminal $c$. This guaranteed that every negative sequence violated at least one deciding predicate. Feature vectors were constructed from all substrings of length at most two, extended with boundary markers, with each feature indicating whether the count of that substring exceeded a given threshold. These features explicitly encode the deciding predicates of $\mathsf{LTT}_{2}$.
@@ Results.
Figure 1 (C) shows the effect of label noise on classification accuracy, and Figure 1 (F) shows the effect of training size. In the noise-free setting, accuracy reached 100%, confirming the theoretical prediction of perfect separability. As noise increased, accuracy decreased, while larger training sizes quickly restored performance toward the ideal. These results provide empirical evidence that threshold-based constraints in $\mathsf{LTT}$ languages are finitely observable and linearly separable when represented with appropriate features.
@@ 6 Experiments on Morphological Data
We represented each affix sequence using two subregular feature classes: Piecewise Testable ($\mathsf{PT}$) predicates, capturing all subsequences up to length $m=2$, and Locally Threshold Testable ($\mathsf{LTT}$) predicates, encoding boundary-aware $k=2$. As in the synthetic experiments, we trained logistic regression classifiers on these feature vectors. Importantly, the classifiers were not constructed directly from the theoretical separating hyperplanes, but were trained from data, demonstrating that linear separability is empirically recoverable by standard learning algorithms.
@@ Results.
On the held-out test set, the classifiers achieved strong performance. Logistic regression with $\mathsf{PT+LTT}$ features typically reached accuracy above 85%, with F1 scores 86%. Confusion matrices in Figure 2 showed that more errors occurred in negative items that were close to positive forms in surface structure. To better understand what the classifier has learned from the real morphological data, we inspected the top-weighted predicates in the logistic regression model (Figure 3). Strikingly, the most influential features correspond to linguistically interpretable affix constraints. For instance, the model strongly weights boundary-sensitive predicates such as the presence of -ly or -ness at the right edge, reflecting the well-attested distributional restrictions of these suffixes. Similarly, certain prefix–suffix interactions emerge as highly discriminative, such as restrictions involving dis- or re- in combination with other affixes. These results demonstrate that the classifier is not only achieving high predictive accuracy, but is also recovering linguistically meaningful generalizations that align with theoretical expectations in morphology.
@@ 7 Conclusion
$\mathsf{SL}_{k}$ $k$-gram presence in $\tilde{x}$ forbidden substrings
$\mathsf{SP}_{k}$ $k$-subsequence presence in $\tilde{x}$ forbidden subsequences
$\mathsf{LT}_{k}$ $k$-grams + (k1)(k{-}1) prefix/suffix local testability
$\mathsf{PT}_{m}$ subsequences of length $\leq m$ Simon’s theorem
$\mathsf{LTT}_{k,\boldsymbol{\tau}}$ thresholded counts for $|g|\leq k$ bounded counting
@@ A.4 Finite Observability of Standard Subregular Classes
Families like $\mathsf{PT}=\bigcup_{m}\mathsf{PT}_{m}$ or $\mathsf{SL}=\bigcup_{k}\mathsf{SL}_{k}$ are directed unions of finitely observable subclasses. Unless parameters are bounded, no single finite predicate set suffices for the whole union; this does not affect applications where the parameters are fixed by design.